# Jacobian Descent for Multi-Objective Optimization

### v2

@@ A.3 Theorem 1
Appendix B Properties of existing aggregators
@@ B.10 Aligned-MTL
Appendix C Experimental settings
@@
C.1 Learning rate selection
@@
C.2 Random reruns and standard error of the mean
To get an idea of confidence in our results, every experiment is performed 8 times on a different seed and a different subset, of size 1024, of the training dataset. The seed used for run $i\in[8]$ is always simply set to $i$. Because each random rerun includes the full learning rate selection method described in Appendix C.1, it is sensible to consider the 8 sets of results as i.i.d. For each point of both the loss curves and the cosine similarity curves, we thus compute the estimated standard error of the mean with the usual formula $\frac{1}{\sqrt{8}}\sqrt{\frac{\sum_{i\in[8]}(v_{i}-\bar{v})^{2}}{8-1}}$, where $v_{i}$ is the value of a point of the curve for random rerun $i$, and $\bar{v}$ is the average value of this point over the 8 runs.
C.3 Model architectures
@@
C.4 Optimizer
For all experiments except those described in Appendix D.3, we always use the basic SGD optimizer of PyTorch, without any regularization or momentum. Here, SGD refers to the PyTorch optimizer that updates the parameters of the model in the opposite direction of the gradient, which, in our case, is replaced by the aggregation of the Jacobian matrix. In the rest of this paper, SGD refers to the whole stochastic gradient descent algorithm. In the experiments of Appendix D.3, we instead use Adam to study its interactions with JD.
C.5 Loss function
@@ C.6 Preprocessing
C.7 Iterations and computational budget
@@
Appendix D Additional experimental results
In this appendix, we provide additional experimental results about IWRM.
D.1 All datasets and all aggregators
Figures 3, 4, 5, 6, 7 and 8 show the full results of the experiments described in Section 5 on SVHN, CIFAR-10, EuroSAT, MNIST, Fashion-MNIST and Kuzushiji-MNIST, respectively. For readability, the results are displayed on three different plots for each dataset. We always show $\mathcal{A}_{\text{UPGrad}}$ and $\mathcal{A}_{\text{Mean}}$ for reference. The exact experimental settings are described in Appendix C.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 3: SVHN results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 4: CIFAR-10 results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 5: EuroSAT results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 6: MNIST results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 7: Fashion-MNIST results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 8: Kuzushiji-MNIST results.
D.2 Varying the batch size
(a) $\text{BS}=4$: Training loss
(b) $\text{BS}=4$: Update similarity to the SGD update
(c) $\text{BS}=16$: Training loss
(d) $\text{BS}=16$: Update similarity to the SGD update
(e) $\text{BS}=64$: Training loss
(f) $\text{BS}=64$: Update similarity to the SGD update
D.3 Compatibility with Adam
Figure 10 gives the results on CIFAR-10 and SVHN when using Adam rather than the SGD optimizer. Concretely, this corresponds to the Adam algorithm in which the gradient is replaced by the aggregation of the Jacobian. The learning rate is still tuned as described in Appendix C.1, but the other hyperparameters of Adam are fixed to the default values of PyTorch, i.e. $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. Because optimization with Adam is faster, the number of epochs for SVHN and CIFAR-10 is reduced to 20 and 15, respectively. While the performance gap is smaller with this optimizer, it is still significant and suggests that our methods are beneficial with other optimizers than the simple SGD. Note that this analysis is fairly superficial. The thorough investigation of the interplay between aggregators and momentum-based optimizers is a compelling future research direction.
(a) SVHN: Training loss
(b) SVHN: Update similarity to the SGD update
(c) CIFAR-10: Training loss
(d) CIFAR-10: Update similarity to the SGD update
Figure 10: Results with the Adam optimizer.
Appendix E Computation time
In this appendix, we compare the computation time of SGD with that of SSJD for all the aggregators that we experimented with. Since we used the same architecture for MNIST, Fashion-MNIST and Kuzushiji-MNIST, we only report the results for one of them. Several factors affect this computation time. First, the batch size affects the number of rows in the Jacobian to aggregate. Increasing the batch size thus requires more GPU memory and the aggregation of a taller matrix. Then, some aggregators, e.g. $\mathcal{A}_{\text{Nash-MTL}}$ and $\mathcal{A}_{\text{MGDA}}$, seem to greatly increase the run time. When the aggregation is the bottleneck, a faster implementation will be necessary to make them usable in practice. Lastly, the current implementation of JD in our library is still fairly inefficient in terms of memory management, which in turn limits how well the GPU can parallelize. Therefore, these results just give a rough indication of the current computation times.
Table 7: Time required in seconds for one epoch of training with SGD and different instances of SSJD, on an NVIDIA L4 GPU. The batch size is always 32.
Method SVHN CIFAR-10 EuroSAT MNIST
SGD 0.79 0.50 0.81 0.47
SSJD + $\mathcal{A}_{\text{Mean}}$ 1.41 1.76 2.93 1.64
SSJD + $\mathcal{A}_{\text{MGDA}}$ 5.50 5.22 6.91 5.22
SSJD + $\mathcal{A}_{\text{DualProj}}$ 1.51 1.88 3.02 1.76
SSJD + $\mathcal{A}_{\text{PCGrad}}$ 2.78 3.13 4.18 3.01
SSJD + $\mathcal{A}_{\text{GradDrop}}$ 1.57 1.90 3.06 1.78
SSJD + $\mathcal{A}_{\text{IMTL-G}}$ 1.48 1.79 2.94 1.69
SSJD + $\mathcal{A}_{\text{CAGrad}}$ 1.93 2.26 3.42 2.17
SSJD + $\mathcal{A}_{\text{RGW}}$ 1.42 1.76 2.89 1.73
SSJD + $\mathcal{A}_{\text{Nash-MTL}}$ 7.88 8.12 9.33 7.91
SSJD + $\mathcal{A}_{\text{Aligned-MTL}}$ 1.53 1.98 2.97 1.71
SSJD + $\mathcal{A}_{\text{UPGrad}}$ 1.80 2.01 3.21 1.90

### v3

@@ B.3 Theorem 1
B.4 Supplementary theoretical results
The proof of Theorem 1 provides some additional insights about the convergence rate as well as some notion of convergence in the non-convex case.
Convergence rate.
Combining the equality ${\bm{f}}^{\ast}={\bm{f}}({\bm{x}}^{\ast})$ and (22), we have
$\frac{1}{T}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t})-{\bm{f}}({\bm{x}}^{\ast})\big{)}\varleq\frac{1}{T}{\bm{1}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{1})-{\bm{f}}({\bm{x}}^{\ast})\big{)}+\frac{\beta\sqrt{m}}{2T}\|{\bm{x}}_{1}-{\bm{x}}^{\ast}\|^{2}$
Furthermore, the Cauchy-Schwartz inequality yields ${\bm{1}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{1})-{\bm{f}}({\bm{x}}^{\ast})\big{)}\varleq\sqrt{m}\cdot\big{\|}{\bm{f}}({\bm{x}}_{1})-{\bm{f}}({\bm{x}}^{\ast})\big{\|}$, so
$\frac{1}{T}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t})-{\bm{f}}({\bm{x}}^{\ast})\big{)}\varleq\frac{\sqrt{m}}{T}\bigg{(}\big{\|}{\bm{f}}({\bm{x}}_{1})-{\bm{f}}({\bm{x}}^{\ast})\big{\|}+\frac{\beta}{2}\|{\bm{x}}_{1}-{\bm{x}}^{\ast}\|^{2}\bigg{)}$ (23)
This hints a convergence rate of order $\mathcal{O}\left(\frac{1}{T}\right)$, whose constant depends on the initial point ${\bm{x}}_{1}$.
Non-convex setting.
The proof of (21) does not require the convexity of the objective function. This bound can be equivalently formulated as
$\frac{1}{T}\sum_{t=1}^{T}\big{\|}J_{t}^{\top}w_{t}\big{\|}^{2}\varleq\frac{2\beta\sqrt{m}}{T}\mathbf{1}^{\top}\big{(}f(x_{1})-f(x^{\ast})\big{)}$ (24)
This shows the convergence of the updates in the non-convex setting, under the smoothness condition of Theorem 1.
Appendix C Properties of existing aggregators
@@ C.10 Aligned-MTL
Appendix D Experimental settings
@@
D.1 Learning rate selection
@@
D.2 Random reruns and standard error of the mean
To get an idea of confidence in our results, every experiment is performed 8 times on a different seed and a different subset, of size 1024, of the training dataset. The seed used for run $i\in[8]$ is always simply set to $i$. Because each random rerun includes the full learning rate selection method described in Appendix D.1, it is sensible to consider the 8 sets of results as i.i.d. For each point of both the loss curves and the cosine similarity curves, we thus compute the estimated standard error of the mean with the usual formula $\frac{1}{\sqrt{8}}\sqrt{\frac{\sum_{i\in[8]}(v_{i}-\bar{v})^{2}}{8-1}}$, where $v_{i}$ is the value of a point of the curve for random rerun $i$, and $\bar{v}$ is the average value of this point over the 8 runs.
D.3 Model architectures
@@
D.4 Optimizer
For all experiments except those described in Appendix A.4, we always use the basic SGD optimizer of PyTorch, without any regularization or momentum. Here, SGD refers to the PyTorch optimizer that updates the parameters of the model in the opposite direction of the gradient, which, in our case, is replaced by the aggregation of the Jacobian matrix. In the rest of this paper, SGD refers to the whole stochastic gradient descent algorithm. In the experiments of Appendix A.4, we instead use Adam to study its interactions with JD.
D.5 Loss function
@@ D.6 Preprocessing
D.7 Iterations and computational budget
@@
Appendix E Computation time and memory usage
E.1 Memory considerations of SSJD for IWRM
The main overhead of SSJD on the IWRM objective comes from having to store the full Jacobian in memory. Remarkably, when we use SGD with ERM, every activation is a tensor whose first dimension is the batch size. Automatic differentiation engines thus have to compute the Jacobian anyway. Since the gradients can be averaged at each layer as soon as they are obtained, the full Jacobian does not have to be stored. In the naive implementation of SSJD, however, storing the Jacobian costs memory, which given the high parallelization ability of GPUs, increases the computational time. With the Gramian-based method proposed in Section 6, only the Gramian, which is typically small, has to be stored: the memory requirement would then be similar to that of SGD.
E.2 Time complexity of the unconflicting projection of gradients
Let $J\in\mathbb{R}^{m\times n}$ be the matrix to aggregate. Apart from computing the Gramian $JJ^{\top}$, applying $\mathcal{A}_{\text{UPGrad}}$ to $J$ requires solving $m$ instances of the quadratic program (5) of Proposition 1. Solvers for such problems of dimension $m$ typically have a computational complexity upper bounded by $\mathcal{O}(m^{4})$ or less in recent implementations (e.g. $\mathcal{O}(m^{3.67}\log m)$). This induces a $\mathcal{O}(m^{5})$ time complexity for extracting the weights of $\mathcal{A}_{\text{UPGrad}}$. Note that solving these $m$ problems in parallel would reduce this complexity to $\mathcal{O}(m^{4})$.
E.3 Empirical computational times
In Table 7, we compare the computation time of SGD with that of SSJD for all the aggregators that we experimented with. Since we used the same architecture for MNIST, Fashion-MNIST and Kuzushiji-MNIST, we only report the results for one of them. Several factors affect this computation time. First, the batch size affects the number of rows in the Jacobian to aggregate. Increasing the batch size thus requires more GPU memory and the aggregation of a taller matrix. Then, some aggregators, e.g. $\mathcal{A}_{\text{Nash-MTL}}$ and $\mathcal{A}_{\text{MGDA}}$, seem to greatly increase the run time. When the aggregation is the bottleneck, a faster implementation will be necessary to make them usable in practice. Lastly, the current implementation of JD in our library is still fairly inefficient in terms of memory management, which in turn limits how well the GPU can parallelize. Also, our implementation of $\mathcal{A}_{\text{UPGrad}}$ does not solve the $m$ quadratic programs in parallel. Therefore, these results just give a rough indication of the current computation times.
Table 7: Time required in seconds for one epoch of training on the ERM objective with SGD and on the IWRM objective with SSJD and different aggregators, on an NVIDIA L4 GPU. The batch size is always 32.
Objective Method SVHN CIFAR-10 EuroSAT MNIST
ERM SGD 0.79 0.50 0.81 0.47
IWRM SSJD–$\mathcal{A}_{\text{Mean}}$ 1.41 1.76 2.93 1.64
IWRM SSJD–$\mathcal{A}_{\text{MGDA}}$ 5.50 5.22 6.91 5.22
IWRM SSJD–$\mathcal{A}_{\text{DualProj}}$ 1.51 1.88 3.02 1.76
IWRM SSJD–$\mathcal{A}_{\text{PCGrad}}$ 2.78 3.13 4.18 3.01
IWRM SSJD–$\mathcal{A}_{\text{GradDrop}}$ 1.57 1.90 3.06 1.78
IWRM SSJD–$\mathcal{A}_{\text{IMTL-G}}$ 1.48 1.79 2.94 1.69
IWRM SSJD–$\mathcal{A}_{\text{CAGrad}}$ 1.93 2.26 3.42 2.17
IWRM SSJD–$\mathcal{A}_{\text{RGW}}$ 1.42 1.76 2.89 1.73
IWRM SSJD–$\mathcal{A}_{\text{Nash-MTL}}$ 7.88 8.12 9.33 7.91
IWRM SSJD–$\mathcal{A}_{\text{Aligned-MTL}}$ 1.53 1.98 2.97 1.71
IWRM SSJD–$\mathcal{A}_{\text{UPGrad}}$ 1.80 2.01 3.21 1.90